Crew 215 Science Report 20Nov2019

[title Science Report – November 20th]

Science Report 20 November 2019
SOL: 10
Crew 215 – Expedition Boomerang Crew
Submitted by Science Officer Andrew Wheeler
Science Dome Operations: The experimenting with gypsum and sand to create a form of
plaster of paris cement continued. 22 cement tiles were created and set aside to dry
for strength testing. The remaining 10 tiles from this programme should be completed
tomorrow. Water liberated from the gypsum has been successfully collected and
approximately 10ml has been bottled. Time permitting tomorrow, an actual gypsum plaster
of paris will be created and the making of splints will be demonstrated.
Green Hab: Green Hab operations have continued with pots filled awaiting the planting
of additional seeds and seedlings. See Green Hab report.
RAM Operations: N/A
EVA Suit Maintenance: Nominal. See Operations Report.
Psycho-social Stressors Monitoring: Data collection continues through the hexoskin(TM)
monitoring vests.
Additional activities: Cleaned airlocks following rain event and attended to three
false alarms of the smoke detectors (caused by water/humidity).

Copyright © The Mars Society. All rights reserved. | Main Site